### v2

# Non-perturbative guiding center model for magnetized plasmas

Abstract
Perturbative guiding center theory adequately describes the slow drift motion of charged particles in the strongly-magnetized regime characteristic of thermal particle populations in various magnetic fusion devices. However, it breaks down for particles with large enough energy. We report on a data-driven method for learning a non-perturbative guiding center model from full-orbit particle simulation data. We show the data-driven model significantly outperforms traditional asymptotic theory in magnetization regimes appropriate for fusion-born $\alpha$-particles in stellarators, thus opening the door to non-perturbative guiding center calculations.
Successful design of a stellarator reactor requires optimizing fusion-born $\alpha$-particle confinement to sustain fusion burn and minimize damage to plasma facing components. Stellarator design efforts therefore require an efficient model for $\alpha$-particle dynamics. State-of-the-art stellarator optimization routines employ the traditional guiding center model Kruskal_1962 ; Littlejohn_1981 ; Littlejohn_1983 ; Littlejohn_1984 ; Cary_2009 ; Burby_loops_JMP , which eliminates the short cyclotron timescale and provides the theoretical basis for advanced confinement concepts such as quasisymmetry Bo2_1983 ; NZ_1988 ; Burby_phase_2013 ; Helander_review_2014 ; BKM_2020 ; Rodriguez_2020 , omnigeneity Hall_1975 ; Helander_review_2014 ; Landreman_2012 ; Parra_2015 , quasi-isodynamicity QI_2023 , and isodrasticity BMN_2023 . This approximation assumes $\epsilon=\rho_{0}/L_{0}\ll 1$, where $\rho_{0}$ denotes a characteristic particle gyroradius and $L_{0}$ denotes a characteristic length scale for the magnetic field $\bm{B}$. When $\epsilon$ is not sufficiently small, traditional guiding center predictions may deviate significantly from those of full-Lorentz-force modeling, leading to imprecise estimates of particle confinement. Because the gyroradius scales linearly with particle speed, energetic $\alpha$-particles may fall in this problematic range even when thermal particles do not. Accuracy of the traditional guiding center model should be critically evaluated in any future stellarator design study.
Figure 1: (Top) Poincaré sections ($z=v_{z}=v_{y}=0,v_{x}>0$) of the full-orbit dynamics of Eq. (1) (black) and level sets of an adiabatic invariant truncation (color) initialized with $v_{x}=\epsilon$ for different values of $\epsilon$. Chaos appears in lobes with smaller field strength due to larger local-$\epsilon$. Lowest-order truncation is used in (a-c), consistent with guiding center models used in practical calculations. $y_{0}$ denotes $y$-coordinate of blue cross. (Bottom) Relative error time traces of three truncations of adiabatic invariant series for the same $\epsilon$ values. Blue crosses denote trajectories underlying (d-f).
Recent numerical studies Assuncao_2023 ; Rodrigues_2024 ; Carbajal_2024  [^1] Ref. Carbajal_2024 is available upon request. Contact: L. Carbajal, leo.carbajal@typeoneenergy.com. find significant deviations in computed $\alpha$-particle loss fractions in various reactor-scale stellarator configurations when using full-orbit vs traditional guiding center particle pushers. These and other calculations indicate that, depending on the stellarator configuration under study, guiding center predictions can either overestimate or underestimate the energetic particle loss fraction computed from full-orbit computations. While some configurations show no discrepancies at all, many lead to differences greater than $60\%$ of the guiding center prediction. Quasisymmetric designs tend to show greater discrepancies than quasi-isodynamic designs. Single trajectory comparisons Assuncao_2023 reveal particles with velocity vectors nearly perpendicular to the magnetic field account for the bulk of these discrepancies. Current designs of some optimized stellarator fusion power plants Hegna_2025 ; Guttenfelder_2025 show good confinement of fusion-born alpha-particles with less than few percent of alpha-particle energy losses. Nevertheless, discrepancy between full-orbit and guiding-center estimates of alpha-particle confinement is still observed Carbajal_2025 , suggesting improvements over guiding center are still necessary for high-fidelity evaluations.
Such deviations may appear surprising because estimating $\epsilon$ for $\alpha$-particles in existing stellarator reactor concepts using nominal minor radius as magnetic scale length $L_{0}$ leads to $\epsilon\sim 2\text{-}3\%$. However, modern stellarator designs often feature highly elongated and curved poloidal cross sections. The local $\epsilon$ in those regions may exceed the nominal value by a factor of  ${\sim}\,5$, leading to a bleaker view of guiding center validity. For example, detailed phase portraits of $\alpha$-particle trajectories in the $2D$ magnetic field $\bm{B}=B(x,y)\,\bm{e}_{z}$, with
$B(x,y)=1+\sum_{i=1}^{2}a_{i}\,\cos(k_{xi}\,x+k_{yi}\,y),$ (1)
$k_{x1}=k_{y2}=3,\quad k_{y1}=k_{x2}=1,\quad a_{1}=a_{2}=0.3,$
reveal significant shifts in particle drift surfaces away from traditional guiding center predictions for $\epsilon\sim 15\%$ (note that $B$ is scaled so $L_{0}=1$). See Fig. 1. Since bounce motion and field line curvature play no role in this field, the quality of traditional guiding center predictions at $\epsilon\sim 15\%$ should only degrade further in a fully $3D$ configuration. Worse still, Fig. 1 reveals that including higher-order perturbative corrections Burby_gc_2013 to the traditional guiding center model, which represents a low-order truncation, offers limited accuracy improvements, if any. This signifies collapse of the optimal truncation order Berry_1991 for the adiabatic invariant asymptotic series. These observations motivate development of a non-perturbative variant of guiding center theory to replace traditional guiding center when it fails.
This Letter describes a non-perturabative guiding center model suitable for $\alpha$-particles in stellarators, runaway electrons in tokamaks Carbajal_2017 ; Carbajal_2020 , and magnetically-confined energetic particles Ogawa_2016 produced by neutral beams Pefferle_2015 , ion cyclotron resonance heating, or fusion reactions Cecconello_2018 more broadly. First we deduce formally-exact non-perturbative guiding center equations of motion assuming a hidden symmetry with associated conserved quantity $\mathcal{J}$. We refer to $\mathcal{J}$ as the non-perturbative adiabatic invariant. These equations are completely determined by first-order derivatives of $\mathcal{J}$ and the magnetic field. They enjoy Hamiltonian structure comparable with that of asymptotic theory Littlejohn_1981 . Then we describe a data-driven method for learning $\mathcal{J}$ from a dataset of full-orbit $\alpha$-particle trajectories. We apply this method to the $\alpha$-particle dynamics shown in Fig. 1 and find the learned non-perturbative guiding center model significantly outperforms the standard guiding center expansion. Our proposed method for learning $\mathcal{J}$ applies on a per-magnetic field basis; changing $\bm{B}$ requires re-training. This makes it well-suited to stellarator design assessment tasks, such as $\alpha$-loss-fraction uncertainty quantification.
In a seminal paper Kruskal_1962 , M. Kruskal showed traditional guiding center theory originates from a hidden perturbative $U(1)$-symmetry of the ordinary differential equations describing single-particle motion, $X=\dot{\bm{x}}\cdot\partial_{\bm{x}}+\dot{\bm{v}}\cdot\partial_{\bm{v}}$, where $\dot{\bm{v}}=\bm{v}\times\bm{B}$ and $\dot{\bm{x}}=\epsilon\,\bm{v}$. Here $U(1)$-symmetry refers to a periodic one-parameter group of phase space transformations that maps solutions of the Lorentz force equations into other solutions. At leading order, the $U(1)$-symmetry is concretely expressed via gyrophase translations. Kruskal dubbed the infinitesimal generator of the perturbative hidden symmetry the roto rate $R$ and showed that its formal expansion in powers of $\epsilon$ is uniquely determined to all orders, with first term given by $R_{0}=\bm{v}\times\bm{b}\cdot\partial_{\bm{v}}$.
Our non-perturbative guiding center model assumes (I) existence of a non-perturbative $U(1)$-symmetry with infinitesimal generator $\mathcal{R}$ such that $\mathcal{R}|_{\epsilon=0}=R_{0}$. We refer to streamlines of $\mathcal{R}$ as $U(1)$-orbits. Here symmetry means the non-perturbative roto rate $\mathcal{R}$ is a Hamiltonian vector field with Hamiltonian function $\mathcal{J}$ that commutes with the kinetic energy $E=|\bm{v}|^{2}/2$ under the Lorentz force Poisson bracket,
$\{F,G\}=\bm{B}\cdot\partial_{\bm{v}}F\times\partial_{\bm{v}}G+\epsilon\,(\partial_{\bm{x}}F\cdot\partial_{\bm{v}}G-\partial_{\bm{x}}G\cdot\partial_{\bm{v}}F),$
i.e. $\{E,\mathcal{J}\}=0$. We refer to $\mathcal{J}$ as the non-perturbative adiabatic invariant.
To ensure topological similarity with the known $U(1)$-symmetry at $\epsilon=0$, we also assume (II) each $U(1)$-orbit generated by $\mathcal{R}$ uniquely and transversally intersects a certain phase space Poincaré section $\Sigma$ defined as follows. Introduce phase space coordinates $(\bm{x},v_{\parallel},v_{\perp},\zeta)$ such that $\bm{v}=v_{\parallel}\,\bm{b}+v_{\perp}\,(\cos\zeta\bm{e}_{1}+\sin\zeta\bm{e}_{2})$. Here $\bm{e}_{1},\bm{e}_{2}$ are unit vector fields chosen to ensure $(\bm{e}_{1},\bm{e}_{2},\bm{b})$ is a right-handed orthonormal frame. Then set $\Sigma=\{(\bm{x},v_{\parallel},v_{\perp},0)\}$. We refer to $\Sigma$ as the guiding center Poincaré section.
Intrinsically, a particle’s guiding center is equal to the $U(1)$-orbit passing through that particle’s phase space location. Guiding center phase space is therefore intrinsically the collection of all $U(1)$-orbits. Constructing a guiding center model, perturbative or non-perturbative, requires equipping this space with coordinates and identifying an evolution law for $U(1)$-orbits in that coordinate system. Traditional guiding center theory painstakingly constructs such coordinates order-by-order in $\epsilon$, resulting in non-unique model equations Parra_Calvo_2014 with exploding complexity that cannot be applied in the non-perturbative regime identified in Fig. 1.
Breaking with tradition, we label $U(1)$-orbits by assigning to each orbit its point of intersection with $\Sigma$, $(\bm{x},v_{\parallel},v_{\perp},\zeta)=(\bm{X},u,r,0)$, or footpoint. In this manner we regard $\Sigma$ as a concrete realization of the abstract space of $U(1)$-orbits. A particle’s footpoint is just the footpoint of the $U(1)$-orbit passing through that particle’s phase space location. We denote the footpoint of a particle at $(\bm{x},v_{\parallel},v_{\perp},\zeta)$ as $\pi(\bm{x},v_{\parallel},v_{\perp},\zeta)=(\bm{X},u,r)$. Note that $\pi(\bm{x},v_{\parallel},v_{\perp},0)=(\bm{x},v_{\parallel},v_{\perp})$, indicating that a particle’s footpoint coincides with its phase space location whenever the particle’s trajectory intersects $\Sigma$, which happens once each gyroperiod. We define the guiding center of any particle as that particle’s footpoint on $\Sigma$.
Figure 2: (a,d) Poincaré sections of the full-orbit dynamics (black) and level sets of the learned $\mathcal{J}_{\Sigma}$ (color) for $\epsilon\in\{0.15,0.27\}$ (cf. Fig. 1 (b,c)). (b,e) Error of $\mathcal{J}_{\Sigma}$ prediction of the dynamics from the full-orbit dynamics over a gyroperiod. (c,f) Error of the order-$j$ truncated dynamics (4) over a gyroperiod with $j=(1,0)$. The error is computed on a uniform $201\times 201$ point grid — a higher resolution than the training data.
As a particle moves through phase space its footpoint traces a curve on $\Sigma$. Because the Lorentz force equations of motion are $U(1)$-invariant by hypothesis that curve is a streamline for a uniquely defined vector field on $\Sigma$ that we call the footpoint flow field, $X_{\Sigma}=\dot{\bm{X}}\cdot\partial_{\bm{X}}+\dot{u}\,\partial_{u}+\dot{r}\,\partial_{r}$. For a proof, see Theorem 1 in supplementary material. The components of the footpoint flow field define our non-perturbative guiding center equations of motion. We identify explicit formulas for these components involving only $\mathcal{J}$, $\bm{B}$, and the unit vectors $\bm{e}_{1},\bm{e}_{2}$ as follows. See Figure 1 in the supplementary material for an illustration of the phase space geometry of footpoint flow.
Any pair of functions $f,g$ on $\Sigma$ determines a pair of $U(1)$-invariant functions $F=f\circ\pi$, $G=g\circ\pi$ on all of phase space by precomposing with the footpoint map $\pi$. Because $\mathcal{R}$ is a Hamiltonian vector field the function $\{F,G\}$ is also $U(1)$-invariant. It follows that the restriction $\{f,g\}_{\Sigma}=\{F,G\}\mid\Sigma$ defines a Poisson bracket on $\Sigma$. This bracket, together with the restricted kinetic energy, $E_{\Sigma}=E\mid\Sigma=u^{2}/2+r^{2}/2$, determines a Hamiltonian system on $\Sigma$. Theorem 2 in supplementary material shows this system coincides with the footpoint flow field. Thus, components of the footpoint flow field $X_{\Sigma}$ may be computed explicitly given an explicit formula for the Poisson bracket $\{f,g\}_{\Sigma}$. This bracket can be expressed entirely in terms of $\mathcal{J}$, $\bm{B}$, and the unit vectors $\bm{e}_{1},\bm{e}_{2}$. The general formula is contained in the proof of Theorem 2 in supplementary material, which includes Ref. Guillemin_Sternberg_1980 . For the $4D$ $(x,y,v_{\perp},\zeta)$-phase space appropriate for magnetic fields of the form $\bm{B}=B(x,y)\,\bm{e}_{z}$ the result is
$\{f,g\}_{\Sigma}$ $=\epsilon\,(\partial_{X}f\,\partial_{r}g-\partial_{r}f\,\partial_{X}g)$
$+\epsilon\,(\partial_{X}\mathcal{J}_{\Sigma}\partial_{r}f-\partial_{r}\mathcal{J}_{\Sigma}\partial_{X}f)\frac{\partial_{r}g+\epsilon\,B^{-1}\,\partial_{Y}g}{\partial_{r}\mathcal{J}_{\Sigma}+\epsilon\,B^{-1}\,\partial_{Y}\mathcal{J}_{\Sigma}}$
$-\epsilon\,(\partial_{X}\mathcal{J}_{\Sigma}\partial_{r}g-\partial_{r}\mathcal{J}_{\Sigma}\partial_{X}g)\frac{\partial_{r}f+\epsilon\,B^{-1}\,\partial_{Y}f}{\partial_{r}\mathcal{J}_{\Sigma}+\epsilon\,B^{-1}\,\partial_{Y}\mathcal{J}_{\Sigma}}.$
Theorem 3 in supplementary material gives explicit formulas for the non-perturbative guiding center equations of motion for general $\bm{B}$. For $\bm{B}=B(x,y)\,\bm{e}_{z}$, the results simplify to $\dot{X}=U/D,\dot{Y}=V/D,\dot{r}=0$, where
$U=-\epsilon^{2}\,\partial_{Y}\mathcal{J}_{\Sigma},\quad V=\epsilon^{2}\,\partial_{X}\mathcal{J}_{\Sigma},$ (2)
$D=(B/r)\,(\partial_{r}\mathcal{J}_{\Sigma}+\epsilon\,B^{-1}\,\partial_{Y}\mathcal{J}_{\Sigma}).$ (3)
Here $\mathcal{J}_{\Sigma}=\mathcal{J}\mid\Sigma$ denotes restriction of the non-perturbative adiabatic invariant to the guiding center Poincaré section. These non-perturbative guiding center evolution laws are exact granted existence of the non-perturbative $U(1)$-symmetry generated by $\mathcal{R}$.
Although formulas (2)-(3) reveal the central role played by the adiabatic invariant in guiding center modeling, they cannot be numerically simulated without an expression for $\mathcal{J}_{\Sigma}$. In the perturbative regime, $\epsilon\ll 1$, truncations of the magnetic moment asymptotic series, such as the second-order result Burby_gc_2013 for $\bm{B}=B(x,y)\,\bm{e}_{z}$,
$\mathcal{J}_{\Sigma}=\frac{r^{2}}{2B}-\epsilon\,\frac{r^{3}\,\partial_{y}B}{2B^{3}}+\epsilon^{2}\frac{r^{4}}{16B^{5}}\left(3\left[\left(\partial_{x}B\right)^{2}+5\left(\partial_{y}B\right)^{2}\right]-B\left(\partial_{x}^{2}B+5\partial_{y}^{2}B\right)\right),$ (4)
provide this missing ingredient. On the other hand, Fig. 1 demonstrates inadequacy of truncated series representations in the non-perturbative regime.
We instead choose to learn $\mathcal{J}_{\Sigma}$ directly from trajectories of the Lorentz force equations for (1) by minimizing the Rayleigh quotient
$\min_{\mathcal{J}_{\Sigma}}\frac{R_{\mathrm{Dyn}}^{2}(\mathcal{J}_{\Sigma})+R_{\mathrm{Inv}}^{2}(\mathcal{J}_{\Sigma})}{\left\lVert\mathcal{J}_{\Sigma}\right\rVert^{2}_{L^{2}(\Omega)}},\quad\mathrm{s.t.}\int_{\Omega}\mathcal{J}_{\Sigma}\,\mathrm{d}^{3}\bm{r}=0,$ (5)
where $\mathcal{J}_{\Sigma}$ is discretized by $39\times 39\times 2$ Fourier by Fourier by Chebyshev modes in $\bm{r}=(x,y,r)\in\Omega=\Omega_{xy}\times[\sqrt{2}-0.01,\sqrt{2}+0.01]$ and $\Omega_{xy}$ is a periodic cell of the magnetic field. Here $\lVert\mathcal{J}_{\Sigma}\rVert_{L^{2}(\Omega)}^{2}=\frac{1}{N}\sum_{i}\mathcal{J}_{\Sigma}(\bm{r}_{i})^{2}$, where the $\bm{r}_{i}$ are sampled from a $79\times 79\times 3$ Fourier by Fourier by Chebyshev-Lobatto quadrature grid on $\Omega$ and $N=51^{2}\times 9$. The first residual
$R_{\mathrm{Dyn}}^{2}=\frac{1}{\hat{N}}\sum_{i}[(D(\hat{\bm{r}}_{i};\mathcal{J}_{\Sigma})\,\dot{X}_{i}-U(\hat{\bm{r}}_{i};\mathcal{J}_{\Sigma}))^{2}+(D(\hat{\bm{r}}_{i};\mathcal{J}_{\Sigma})\,\dot{Y}_{i}^{n}-V(\hat{\bm{r}}_{i};\mathcal{J}_{\Sigma}))^{2}]$
penalizes differences between predictions from $\mathcal{J}_{\Sigma}$ and full-orbit trajectories on the Poincaré section. A similar objective function appeared previously in Messenger_2024 , where a parametric averaged Hamiltonian appeared in place of our parametric $\mathcal{J}_{\Sigma}$. For the sum we initially sample $1000$ points $\hat{\bm{r}}_{i}$ from a low-discrepancy sequence on $\Omega$. For the $\hat{N}$ points that correspond to integrable trajectories, we estimate footpoint time derivatives $(\dot{X}_{i},\dot{Y}_{i})$ using a technique based on Ruth:2024 . (See supplementary materials.) The residual $R_{\mathrm{Dyn}}^{2}$ ignores non-integrable trajectories. The second residual is given by $R_{\mathrm{Inv}}^{2}=\frac{1}{N}\sum_{i}(\mathcal{J}_{\Sigma}(\bm{r}_{i})-\mathcal{J}_{\Sigma}(F(\bm{r}_{i})))^{2}$, where $F:\Omega\to\Omega$ denotes the full-orbit Poincaré map on $\Sigma$. It penalizes non-invariance of $\mathcal{J}_{\Sigma}$ Ruth:2023 . This, along with the oversampling in $\bm{r}_{i}$, smoothes $\mathcal{J}_{\Sigma}$ in chaotic regions. The quotient can be minimized via a single generalized eigenvalue problem because both residuals in (5) are quadratic in $\mathcal{J}_{\Sigma}$. Computing the $\hat{N}$ samples of $({X},{Y})$ dominates computational time. An average trajectory computation requires around $1000$ gyroperiods, so the method’s total cost is on the order of the time to simulate $1000$ gyroperiods for $1000$ particles using a full-orbit integrator.
Many other methods compute approximate invariant structures, including Lagrangian coherent structures hadjighasem2017 ; froyland2014 , flux surfaces dewar-qfm-maps ; dewar-qfm-fields ; Ruth:2023 , and invariant tori more generally haro-parameterization . However, these methods do not distinguish between the true adiabatic invariant $\mathcal{J}$ associated with continuation of Kruskal’s perturbative $U(1)$-symmetry and arbitrary invariants of the form $I(E,\mathcal{J})$. Our method accounts for this ambiguity because replacing $\mathcal{J}$ with $I(E,\mathcal{J})$ in the non-perturbative equations of motion (2) generally changes predicted footpoint trajectories, and thereby increases the residual $R_{\mathrm{Dyn}}^{2}(\mathcal{J}_{\Sigma})$. In contrast to general constants of motion, the learned non-perturbative invariant $\mathcal{J}$ completely determines the dynamics.
Fig. 2 compares predictions using learned $\mathcal{J}_{\Sigma}$ to full-orbit simulation data previously shown in Fig. 1 at $\epsilon\in\{0.15,0.27\}$. The global phase portraits in Fig. 2 (a,d) reveal significant improvements over the asymptotics shown in Fig. 1 (b,c). Let $F_{\mathcal{J}}$ denote the approximate Poincaré map obtained by evolving Eqs. (2)-(3) over a gyroperiod using learned $\mathcal{J}$. Let $F_{j}$ denote the corresponding map using order-$j$ truncation of (4). Figure 2 (b,e) compares the log-absolute errors, $\log_{10}|F_{\mathcal{J}}-F|$ and $\log_{10}|F_{j}-F|$, with $j=1$ for $\epsilon=0.15$ and $j=0$ for $\epsilon=0.27$ (larger $j$ leads to zero denominators $D$ in both cases). For both $\epsilon$-values, Poincaré map dynamics predicted using learned $\mathcal{J}_{\Sigma}$ outperforms traditional asymptotics by orders of magnitude over most of the phase portrait.
Discussion– Advanced stellarator design principles like quasisymmetry and omnigeneity aim to improve confinement by imbuing the traditional guiding center model with desirable dynamical features, such as symmetry or small radial flux. Reliance on traditional guiding center limits applicability of these principles to the regime $\epsilon\ll 1$. New confinement principles targeting comparable dynamical improvements for the non-perturbative guiding center model promise to extend the benefits of optimized confinement to larger-$\epsilon$ particles, such as $\alpha$-particles in stellarators that break the traditional guiding center model.
Assessing a candidate stellarator reactor design merits sensitivity analysis of $\alpha$-particle loss fraction over an energy confinement time. Assessing sensitivity to perturbations of the $\alpha$-particle distribution function requires either high-dimensional adjoint solves or large Monte-Carlo (MC) sampling of $\alpha$-particle trajectories from various perturbed distributions. Because each MC sample evolves in the same field configuration, computational overhead associated with training the non-perturbative guiding center model falls well below the cost of performing all required Monte-Carlo sampling using a full-orbit approach. Thus, in device designs where guiding center fails for $\alpha$-particles, accuracy and timestepping efficiency motivate trajectory sampling using the non-perturbative guiding center model over traditional guiding center or full-orbit models.
Direct optimization of $\alpha$-particle confinement albert_accelerated_2020 ; albert_alpha_2023 ; bindel_direct_2023 entails repeated ensemble trajectory simulations during an inner field-optimization loop. Computational feasibility of this technique currently requires traditional guiding center for trajectory calculations, limiting its applicability to particles with $\epsilon\ll 1$. Deploying a trained non-perturbative guiding center model in place of traditional guiding center promises to extend applicability of direct fast-particle confinement optimization to particles with non-perturbative $\epsilon$. However, the training method proposed here falls short for inner-loop applications because it must be repeated whenever the field changes, thereby encurring unacceptable computational overhead comparable to full-orbit trajectory generation. Overcoming this barrier requires solving the operator learning problem LuLu_2021 of mapping $\bm{B}$ to $\mathcal{J}$. Recent advances in machine learning, especially in foundation models Subramanian_2023 and sparse regression Brunton_2016 ; Messenger_2021 , stand poised to tackle this challenge.
[…]
Acknowledgements– This material is based on work supported by the U.S. Department of Energy, Office of Science, Office of Advanced Scientific Computing Research, as a part of the Mathematical Multifaceted Integrated Capability Centers program, under Award Number DE-SC0023164. It was also supported by U.S. Department of Energy grant # DE-FG02-04ER54742, as well as the Los Alamos National Laboratory LDRD program under Project No. 20240858PRD.
I Supplementary material
I.1 Dynamical systems underpinning
This Section provides statements and proofs of some basic results in the theory of ordinary differential equations (ODEs) with $U(1)$-symmetry and Hamiltonian structure. It also shows how these results generalize the theory presented in the main text from magnetic fields of the special form $\bm{B}=B(x,y)\,\bm{e}_{z}$ to general non-vanishing magnetic fields. Throughout, $\dot{z}=X(z)$ denotes a system of first-order ODEs on the phase space $Z\ni z$. Note that $X$ may be understood as a vector field on $Z$ that encodes the ODE. We make no distinction between the ODE system and the vector field $X$. We will always use the symbol $F_{t}=\exp(t\,X):Z\rightarrow Z$ to denote the time-$t$ flow map for $X$. This discussion assumes smoothness of the various geometric objects that appear.
As is standard, the symbol $U(1)$ denotes the group of complex numbers with unit modulus $e^{i\theta}$. We identify this group with the set of real numbers $\theta$ modulo $2\pi$. To formalize the notion of $U(1)$-symmetry we refer to $U(1)$-actions. A $U(1)$-action on $Z$ is a family of mappings $\Phi_{\theta}:Z\rightarrow Z$, parameterized by $\theta\in U(1)$, such that $\Phi_{0}=\Phi_{2\pi}=\text{id}_{Z}$ and $\Phi_{\theta_{1}+\theta_{2}}=\Phi_{\theta_{1}}\circ\Phi_{\theta_{2}}$ for each $\theta_{1},\theta_{2}\in U(1)$.
We think of a $U(1)$-action as a collection of generalized rotations. Given $z\in Z$ the $U(1)$-orbit containing $z$ is the set $\mathcal{O}_{z}=\{\Phi_{\theta}(z)\mid\theta\in U(1)\}$ comprising all possible rotations of $z$. The vector field $R(z)=(\partial_{\theta}\Phi_{\theta}(z))|_{\theta=0}$ is the infinitesimal generator of the $U(1)$-action, which is tangent to the collection of $U(1)$-orbits. We say that $X$ is $U(1)$-invariant with respect to a $U(1)$-action $\Phi_{\theta}$ if $F_{t}\circ\Phi_{\theta}=\Phi_{\theta}\circ F_{t}$ for every $\theta\in U(1)$ and $t\in\mathbb{R}$. If the $U(1)$-action is contextually clear, we will simply say $X$ is $U(1)$-invariant.
Fix a $U(1)$-action $\Phi_{\theta}$ and suppose that $X$ is $U(1)$-invariant. Assume there is a global Poincaré section $\Sigma\subset Z$ for $\Phi_{\theta}$ Then, by definition of global Poincaré sections, $\Sigma$ is a hypersurface in $Z$ and each $U(1)$-orbit intersects $\Sigma$ uniquely and transversally. Let $\pi:Z\rightarrow\Sigma$ denote the mapping that sends $z\in Z$ to the unique point of intersection between $\Sigma$ and $\mathcal{O}_{z}$. The main text refers to $\pi$ as the footpoint map. A basic result referred to in the main text shows that $\pi$ maps the ODE system $X$ on $Z$ to another ODE system $X_{\Sigma}$ (i.e. a vector field) on $\Sigma$. The main text refers to $X_{\Sigma}$ as the footpoint flow field.
Theorem 1.
There is a unique vector field $X_{\Sigma}$ on $\Sigma$ such that, for every streamline $z(t)$ of $X$, $\sigma(t)=\pi(z(t))$ is a streamline of $X_{\Sigma}$.
Proof.
Commutativity of $F_{t}$ and $\Phi_{\theta}$ implies there is a $1$-parameter family of mappings $f_{t}:\Sigma\rightarrow\Sigma$, $t\in\mathbb{R}$, such that $\pi\circ F_{t}=f_{t}\circ\pi$. To see this, let $z_{0}\in\mathcal{O}_{\sigma}$ be an arbitrary point in the $U(1)$-orbit containing $\sigma\in\Sigma$. The image point $z(t)=F_{t}(z_{0})$ is contained in the $U(1)$-orbit $\mathcal{O}_{\pi(z(t))}$. If $z_{0}^{\prime}\in\mathcal{O}_{\sigma}$ is any other point in the $U(1)$-orbit containing $\sigma$ then there is a $\theta\in U(1)$ such that $z_{0}^{\prime}=\Phi_{\theta}(z_{0})$. By commutativity, the new image point $z^{\prime}(t)=F_{t}(z_{0}^{\prime})$ is related to the previous one according to $z^{\prime}(t)=F_{t}(\Phi_{\theta}(z_{0}))=\Phi_{\theta}(F_{t}(z_{0}))=\Phi_{\theta}(z(t))$. Thus, $z^{\prime}(t)$ and $z(t)$ lie on a common $U(1)$-orbit. The projected image $\pi(F_{t}(\mathcal{O}_{\sigma}))$ is therefore the singleton set $\{\sigma(t)\}$, where $\sigma(t)=\pi(F_{t}(z_{0}))$, for any $z_{0}\in\mathcal{O}_{\sigma}$. We define $f_{t}(\sigma)=\sigma(t)$. Let $z_{0}\in Z$ be any point in phase space and set $\sigma=\pi(z_{0})$. By definition of $f_{t}$, we have
$f_{t}(\pi(z_{0}))=f_{t}(\sigma)=\sigma(t)=\pi(F_{t}(z_{0})),$
as claimed.
The commuting property $\pi\circ F_{t}=f_{t}\circ\pi$ implies the $1$-parameter family of mappings $f_{t}:\Sigma\rightarrow\Sigma$ satisfies the flow property $\forall t_{1},t_{2}\in\mathbb{R},\quad f_{t_{1}+t_{2}}=f_{t_{1}}\circ f_{t_{2}}$. Indeed, if $\sigma\in\Sigma$ there is some $z\in Z$ with $\pi(z)=\sigma$, which implies
$f_{t_{1}+t_{2}}(\sigma)=f_{t_{1}+t_{2}}(\pi(z))$ $=\pi(F_{t_{1}+t_{2}}(z))=\pi(F_{t_{1}}(F_{t_{2}}(z)))$
$=f_{t_{1}}(\pi(F_{t_{2}}(z)))=f_{t_{1}}(f_{t_{2}}(\pi(z)))=f_{t_{1}}(f_{t_{2}}(\sigma)).$
Therefore $f_{t}$ is the flow map for a vector field $X_{\Sigma}$ on $\Sigma$.
If $z(t)$ is any streamline for $X$ then $\sigma(t)=\pi(z(t))$ is a streamline for $X_{\Sigma}$. To see this first let $z_{0}=z(0)$, $\sigma_{0}=\sigma(0)$, and observe that $\sigma(t)=\pi(F_{t}(z_{0}))=f_{t}(\sigma_{0})$. Since $f_{t}$ is the flow map for $X_{\Sigma}$ it follows that $\sigma(t)$ is a streamline for $X_{\Sigma}$.
Suppose that $Y_{\Sigma}$ were a second vector field on $\Sigma$ sharing the previous property with $X_{\Sigma}$. For $\sigma_{0}\in\Sigma$ let $z_{0}$ be any point in the $U(1)$-orbit containing $\sigma_{0}$. There is a unique $X$-streamline $z(t)$ with $z(0)=z_{0}$. Moreover $\sigma(t)=\pi(z(t))$ is a streamline for both $X_{\Sigma}$ and $Y_{\Sigma}$. In particular,
$Y_{\Sigma}(\sigma_{0})=Y_{\Sigma}(\sigma(0))=\frac{d}{dt}\bigg{|}_{0}\sigma(t)=X_{\Sigma}(\sigma(0))=X_{\Sigma}(\sigma_{0}),$
which implies $X_{\Sigma}=Y_{\Sigma}$. ∎
Remark 1.
This result has an interesting interpretation when $\Sigma$ is also a Poincaré section for $X$, the ODE system of interest, as happens in the case of charged particles moving in a strong magnetic field. Since $\Sigma$ is a Poincaré section for $X$ there is a well-defined poincaré map and associated discrete-time dynamics for $X$ on $\Sigma$. It is generally interesting to inquire as to whether there is a continuous-time dynamical system on $\Sigma$ that provides continuous-time interpolation of discrete-time Poincaré map dynamics. Theorem 1 says that, owing to the presence of $U(1)$-symmetry with a common Poincaré section, there is indeed such a dynamical system on $\Sigma$ – that defined by the footpoint flow field $X_{\Sigma}$.
Remark 2.
Fig. 3 gives a visual representation of Theorem 1.
Figure 3: An illustration of a trajectory described by Theorem 1. The full orbit trajectory is given in blue and the corresponding footpoint trajectory is in black, traveling on the gray Poincaré section $\Sigma$. An example of a $U(1)$-symmetry orbit is plotted in red. Although only a single $U(1)$-orbit is shown, there is such an orbit containing $z(t)$ for all times $t$.
Without further information about the ODE system $X$ finding its footpoint flow field requires detailed knowledge of the $U(1)$-orbits. These orbits are known when the underlying $U(1)$-symmetry is known in advance, but not when dealing with “hidden” symmetries, as in non-perturbative guiding center modeling. Fortunately, matters simplify considerably when $X$ is a Hamiltonian system like the Lorentz force Law. The following Theorem shows that in the Hamiltonian setting the footpoint flow field is completely determined by $\Sigma$, the Poisson bracket for $X$, the Hamiltonian for $X$, and the conserved quantity $J$ associated with the $U(1)$-action $\Phi_{\theta}$; detailed knowledge of the $U(1)$-orbits is not required.
Theorem 2.
[…]
  • the infinitesimal generator $R$ for the $U(1)$-action is a Hamiltonian vector field with Hamiltonian $J$,
then the following is true.
  1. 1.
    The hypersurface $\Sigma$ has a natural Poisson bracket {,}$\{\cdot,\cdot\}_{\Sigma}$ with $J_{\Sigma}=J\mid\Sigma$ as a Casimir invariant. If $Z$ is equipped with coordinates $(\sigma^{i},\zeta)$, where $\zeta$ is an angular coordinate, and $\Sigma=\{\zeta=0\}$ then the $\sigma^{i}$ parameterize $\Sigma$. Moreover the Poisson bracket $\{f,g\}_{\Sigma}$ between functions $f=f(\sigma)$, and $g=g(\sigma)$ on $\Sigma$ is given explicitly by
    $\{f,g\}_{\Sigma}$ $=\partial_{\sigma}f^{T}\mathbb{J}_{\Sigma}\partial_{\sigma}g+\partial_{\sigma}J_{\Sigma}^{T}\mathbb{J}_{\Sigma}\,\partial_{\sigma}f\,\frac{N_{\Sigma}^{T}\partial_{\sigma}g}{N_{\Sigma}^{T}\partial_{\sigma}J_{\Sigma}}-\partial_{\sigma}J_{\Sigma}^{T}\mathbb{J}_{\Sigma}\,\partial_{\sigma}g\,\frac{N_{\Sigma}^{T}\partial_{\sigma}f}{N_{\Sigma}^{T}\partial_{\sigma}J_{\Sigma}},$ (6)
    where the matrix $\mathbb{J}_{\Sigma}$ has components $\mathbb{J}_{\Sigma}^{ij}=\{\sigma^{i},\sigma^{j}\}(\sigma,0)$, and the column vector $N$ has components $N^{i}_{\Sigma}=\{\sigma^{i},\zeta\}(\sigma,0)$.
  2. 2.
    The footpoint flow field $X_{\Sigma}$ on $\Sigma$ is a Hamiltonian vector field with respect to the Poisson bracket {,}$\{\cdot,\cdot\}_{\Sigma}$ and the Hamiltonian function $H_{\Sigma}=H\mid\Sigma$.
Proof.
[…]
$\mathcal{L}_{R}\{F,G\}=\{\{F,G\},J\}=\{\{F,J\},G\}+\{F,\{G,J\}\}=\{\mathcal{L}_{R}F,G\}+\{F,\mathcal{L}_{R}G\},$
[…]
Let $\iota_{\Sigma}:\Sigma\rightarrow Z$ denote the inclusion map for the Poincaré section $\Sigma$. Given smooth functions $f,g:\Sigma\rightarrow\mathbb{R}$ define their bracket according to
$\{f,g\}_{\Sigma}=\iota_{\Sigma}^{*}\{\pi^{*}f,\pi^{*}g\},$
where $\pi:Z\rightarrow\Sigma$ denotes the footpoint map. This bracket is clearly bilinear and skew-symmetric. It satisfies the Leibniz property because
$\{f,gh\}_{\Sigma}$ $=\iota_{\Sigma}^{*}\bigg{(}\{\pi^{*}f,\pi^{*}g\}\,\pi^{*}h+\{\pi^{*}f,\pi^{*}h\}\,\pi^{*}g\bigg{)}$
=$=\{f,g\}_{\Sigma}\,\iota_{\Sigma}^{*}\pi^{*}h+\{f,h\}_{\Sigma}\,\iota_{\Sigma}^{*}\pi^{*}g$
=$=\{f,g\}_{\Sigma}\,h+\{f,h\}_{\Sigma}\,g,$
[…]
$\{\{f,g\}_{\Sigma},h\}_{\Sigma}$ $=\iota_{\Sigma}^{*}\{\pi^{*}\iota_{\Sigma}^{*}\{\pi^{*}f,\pi^{*}g\},\pi^{*}h\}=\iota_{\Sigma}^{*}\{\{\pi^{*}f,\pi^{*}g\},\pi^{*}h\},$
[…]
$\{f,J_{\Sigma}\}_{\Sigma}=\iota_{\Sigma}^{*}\{\pi^{*}f,\pi^{*}J_{\Sigma}\}=\iota_{\Sigma}^{*}\{\pi^{*}f,J\}=\iota_{\Sigma}^{*}(\mathcal{L}_{R}\pi^{*}f)=0,$
for each smooth function $f$ on $\Sigma$.
The footpoint map $\pi:Z\rightarrow\Sigma$ is a Poisson map between $(Z,\{\cdot,\cdot\})$ and $(\Sigma,\{\cdot,\cdot\}_{\Sigma})$ because
$\pi^{*}\{f,g\}_{\Sigma}=\pi^{*}\iota_{\Sigma}^{*}\{\pi^{*}f,\pi^{*}g\}=\{\pi^{*}f,\pi^{*}g\}.$
The Hamiltonian $H$ is $U(1)$-invariant by hypothesis and therefore determined by its values on the Poincaré section, i.e. by the restriction $H_{\Sigma}=H\mid\Sigma$, in the sense that $\pi^{*}H_{\Sigma}=H$. It now follows from Guillemin-Sternberg collectivization Guillemin_Sternberg_1980 that if $z(t)$ is any solution of Hamilton’s equations on $Z$ with Hamiltonian $H=\pi^{*}H_{\Sigma}$ then the corresponding footpoint trajectory $\sigma(t)=\pi(z(t))$ is a solution of Hamilton’s equations on $\Sigma$ with Hamiltonian $H_{\Sigma}$. In other words, the footpoint trajectories are streamlines for the Hamiltonian vector field on $\Sigma$ with Hamiltonian $H_{\Sigma}$. But, by Theorem 1, the only vector field on $\Sigma$ that has footpoint trajectories as streamlines is the footpoint flow field $X_{\Sigma}$. This establishes the second part of the theorem.
To complete the proof, we must now confirm the formula (6) for the Poisson bracket {,}$\{\cdot,\cdot\}_{\Sigma}$ in the coordinates $(\sigma^{i},\zeta)$. Without loss of generality, suppose the index $i$ takes values in $\{1,\dots,d\}$, for some integer $d$. There must be a skew-symmetric matrix $\mathbb{J}^{ij}(\sigma,\zeta)$ and a column vector $N^{i}(\sigma,\zeta)$ such that
$\{F,G\}$ $=\begin{pmatrix}\partial_{\sigma^{1}}F&\dots&\partial_{\sigma^{d}}F&\partial_{\zeta}F\end{pmatrix}\begin{pmatrix}\mathbb{J}&N\\ -N^{T}&0\end{pmatrix}\begin{pmatrix}\partial_{\sigma^{1}}G\\ \vdots\\ \partial_{\sigma^{d}}G\\ \partial_{\zeta}G\end{pmatrix}$
$=\partial_{\sigma}F^{T}\mathbb{J}\partial_{\sigma}G+\partial_{\zeta}G\,N^{T}\partial_{\sigma}F-\partial_{\zeta}F\,N^{T}\partial_{\sigma}G.$
Setting $F=\sigma^{i}$ and $G=\sigma^{j}$ shows that the components of $\mathbb{J}$ are given by
$\mathbb{J}^{ij}=\{\sigma^{i},\sigma^{j}\},$
while setting $F=\sigma^{i}$ and $G=\zeta$ shows that the components of $N$ are given by
$N^{i}=\{\sigma^{i},\zeta\}.$
The inclusion map $\iota_{\Sigma}:\Sigma\rightarrow Z$ is given by $\sigma\mapsto(\sigma,0)$ in these coordinates. The Poisson bracket {,}$\{\cdot,\cdot\}_{\Sigma}$ is therefore given by
$\{f,g\}_{\Sigma}=\partial_{\sigma}f^{T}\mathbb{J}_{\Sigma}\partial_{\sigma}g+\partial_{\zeta}(\pi^{*}g)(\sigma,0)\,N_{\Sigma}^{T}\partial_{\sigma}f-\partial_{\zeta}(\pi^{*}f)(\sigma,0)\,N_{\Sigma}^{T}\partial_{\sigma}g,$ (7)
where $\mathbb{J}_{\Sigma}=\mathbb{J}(\sigma,0)$ and $N_{\Sigma}=N(\sigma,0)$. This would be an explicit formula for $\{f,g\}_{\Sigma}$ were it not for the appearance of the $\zeta$-derivatives $\partial_{\zeta}(\pi^{*}f)(\sigma,0)$ and $\partial_{\zeta}(\pi^{*}g)(\sigma,0)$. These $\zeta$-derivatives can be expressed in terms of derivatives of $f,g$ as follows. By definition of the footpoint map $\pi$, the functions $\pi^{*}f,\pi^{*}g$ are each constant along $U(1)$-orbits. Equivalently, $\mathcal{L}_{R}\pi^{*}f=\mathcal{L}_{R}\pi^{*}g=0$. Writing the infinitesimal generator in components as $R=R^{\zeta}\,\partial_{\zeta}+\bm{R}^{i}\,\partial_{\sigma^{i}}$ therefore implies $R^{\zeta}\,\partial_{\zeta}(\pi^{*}f)+\bm{R}^{T}\,\partial_{\sigma}(\pi^{*}f)=0$, or
$\partial_{\zeta}(\pi^{*}f)=-\frac{\bm{R}^{T}\,\partial_{\sigma}(\pi^{*}f)}{R^{\zeta}},\quad\partial_{\zeta}(\pi^{*}g)=-\frac{\bm{R}^{T}\,\partial_{\sigma}(\pi^{*}g)}{R^{\zeta}}.$
These expressions can be simplified further using the fact that $J$ is the Hamiltonian for $R$. In particular, the components of $R$ must be given in terms of derivatives of $J$ according to
$R^{\zeta}=\{\zeta,J\}=-N^{T}\partial_{\sigma}J,\quad\bm{R}=\{\sigma,J\}=\mathbb{J}\partial_{\sigma}J+N\,\partial_{\zeta}J.$
The $\zeta$-derivatives are therefore
$\partial_{\zeta}(\pi^{*}f)=\frac{(\mathbb{J}\partial_{\sigma}J+N\,\partial_{\zeta}J)^{T}\,\partial_{\sigma}(\pi^{*}f)}{N^{T}\partial_{\sigma}J},\quad\partial_{\zeta}(\pi^{*}g)=\frac{(\mathbb{J}\partial_{\sigma}J+N\,\partial_{\zeta}J)^{T}\,\partial_{\sigma}(\pi^{*}g)}{N^{T}\partial_{\sigma}J},$
and the Poisson bracket (7) becomes
$\{f,g\}_{\Sigma}$ $=\partial_{\sigma}f^{T}\mathbb{J}_{\Sigma}\partial_{\sigma}g$
$+\frac{(\mathbb{J}_{\Sigma}\partial_{\sigma}J_{\Sigma}+N_{\Sigma}\,\partial_{\zeta}J(\sigma,0))^{T}\,\partial_{\sigma}g}{N_{\Sigma}^{T}\partial_{\sigma}J_{\Sigma}}\,N_{\Sigma}^{T}\partial_{\sigma}f$
$-\frac{(\mathbb{J}_{\Sigma}\partial_{\sigma}J_{\Sigma}+N_{\Sigma}\,\partial_{\zeta}J(\sigma,0))^{T}\,\partial_{\sigma}f}{N_{\Sigma}^{T}\partial_{\sigma}J_{\Sigma}}\,N_{\Sigma}^{T}\partial_{\sigma}g$
$=\partial_{\sigma}f^{T}\mathbb{J}_{\Sigma}\partial_{\sigma}g+\frac{(\mathbb{J}_{\Sigma}\partial_{\sigma}J_{\Sigma})^{T}\,\partial_{\sigma}g}{N_{\Sigma}^{T}\partial_{\sigma}J_{\Sigma}}\,N_{\Sigma}^{T}\partial_{\sigma}f-\frac{(\mathbb{J}_{\Sigma}\partial_{\sigma}J_{\Sigma})^{T}\,\partial_{\sigma}f}{N_{\Sigma}^{T}\partial_{\sigma}J_{\Sigma}}\,N_{\Sigma}^{T}\partial_{\sigma}g$
$=\partial_{\sigma}f^{T}\mathbb{J}_{\Sigma}\partial_{\sigma}g+\partial_{\sigma}J_{\Sigma}^{T}\mathbb{J}_{\Sigma}\,\partial_{\sigma}f\,\frac{N_{\Sigma}^{T}\partial_{\sigma}g}{N_{\Sigma}^{T}\partial_{\sigma}J_{\Sigma}}-\partial_{\sigma}J_{\Sigma}^{T}\mathbb{J}_{\Sigma}\,\partial_{\sigma}g\,\frac{N_{\Sigma}^{T}\partial_{\sigma}f}{N_{\Sigma}^{T}\partial_{\sigma}J_{\Sigma}}.$
Remark 3.
For the purposes of non-perturbative guiding center modeling, $\Sigma$; the Poisson bracket for $X$; and the Hamiltonian for $X$; are each known explicitly in advance. Theorem 2 therefore implies that the conserved quantity $\mathcal{J}$ associated with the “hidden” $U(1)$-symmetry of the Lorentz force Law is all that is needed to find an explicit formula for the footpoint flow field, and therefore the non-perturbatibe guiding center model. The underlying mechanism that enables this remarkable simplification is Noether’s theorem. Finding a formula for the footpoint flow field in general requires detailed knowledge of the underlying $U(1)$-symmetry. On the other hand, Noether’s theorem implies that any Hamiltonian $U(1)$-symmetry is completely determined by its corresponding conservation law. Thus, in the Hamiltonian setting complete knowledge of the Noether conserved quantity implies complete knowledge of the corresponding $U(1)$-symmetry and therefore complete knowledge of the footpoint flow field.
When Theorem 2 is applied to the Lorentz force Law written in dimensionless variables as $\dot{\bm{v}}=\bm{v}\times\bm{B}(\bm{x})$, $\dot{\bm{x}}=\epsilon\,\bm{v}$ it leads to the following characterization of the non-perturbative guiding center equations of motion in general non-vanishing magnetic fields. This result shows explicitly how to remove the assumption $\bm{B}=B(x,y)\,\bm{e}_{z}$ used in the main text. It also recovers the main text’s non-perturbative guiding center equations of motion when $\bm{B}$ is translation-invariant along $z$.
Theorem 3.
Parameterize the guiding center Poincaré section $\Sigma$ using coordinates $\bm{\mathsf{X}}=\bm{x}$, $u=v_{\parallel}$, $r=v_{\perp}$. Let $J_{\Sigma}$ denote the restriction of the non-perturbative action integral to $\Sigma$. The footpoint flow field $X_{\Sigma}=\dot{\bm{\mathsf{X}}}\cdot\partial_{\bm{\mathsf{X}}}+\dot{u}\,\partial_{u}+\dot{r}\,\partial_{r}$ on $\Sigma$ for the Lorentz force system is given explicitly by
$D_{\epsilon}\,\dot{\bm{\mathsf{X}}}\cdot\bm{b}$ $=\epsilon\,D_{\epsilon}\,u-{\epsilon\,\partial_{u}J_{\Sigma}\bigg{(}|\bm{B}|+\epsilon\,{(u\,\bm{b}+r\,\bm{e}_{1})\cdot[u\,r^{-1}\,\nabla\bm{b}\cdot\bm{e}_{2}+\nabla\bm{e}_{1}\cdot\bm{e}_{2}]}\bigg{)}}$ (8)
$D_{\epsilon}\,\dot{\bm{\mathsf{X}}}\cdot\bm{e}_{1}$ $=\epsilon\,D_{\epsilon}\,r-\epsilon\,{\partial_{r}J_{\Sigma}\bigg{(}|\bm{B}|+\epsilon\,{(u\,\bm{b}+r\,\bm{e}_{1})\cdot[u\,r^{-1\,}\nabla\bm{b}\cdot\bm{e}_{2}+\nabla\bm{e}_{1}\cdot\bm{e}_{2}]}\bigg{)}}$ (9)
$D_{\epsilon}\,\dot{\bm{\mathsf{X}}}\cdot\bm{e}_{2}$ $=-\epsilon^{2}\,{r^{-1}(u\,\bm{b}+r\,\bm{e}_{1})\cdot\partial_{\bm{\mathsf{X}}}J_{\Sigma}+\epsilon^{2}\,r^{-1}(u\,\bm{b}+r\,\bm{e}_{1})\cdot(\nabla\bm{b}\cdot\bm{e}_{1})(r\,\partial_{u}J_{\Sigma}-u\,\partial_{r}J_{\Sigma})}$ (10)
$D_{\epsilon}\,\dot{u}$ $=\epsilon\,\bigg{(}\bm{B}+\epsilon\,(\bm{b}\cdot\nabla\times\bm{b})[u\,\bm{b}+r\,\bm{e}_{1}]$
$\hskip 30.00005pt+\epsilon\,\bm{b}\times([u\,\bm{b}+r\,\bm{e}_{1}]\cdot\nabla\bm{b})-\epsilon\,r\,\bm{e}_{2}\times(\nabla\bm{e}_{1}\cdot\bm{e}_{2})\bigg{)}\cdot\partial_{\bm{\mathsf{X}}}J_{\Sigma}$ (11)
$D_{\epsilon}\,\dot{r}$ $=-\epsilon\,u\,r^{-1}\,\bigg{(}\bm{B}+\epsilon\,(\bm{b}\cdot\nabla\times\bm{b})[u\,\bm{b}+r\,\bm{e}_{1}]$
$\hskip 60.00009pt+\epsilon\,\bm{b}\times([u\,\bm{b}+r\,\bm{e}_{1}]\cdot\nabla\bm{b})-\epsilon\,r\,\bm{e}_{2}\times(\nabla\bm{e}_{1}\cdot\bm{e}_{2})\bigg{)}\cdot\partial_{\bm{\mathsf{X}}}J_{\Sigma},$ (12)
where the denominator $D$ is given by
$D_{\epsilon}$ $=\epsilon\,r^{-1}\bm{e}_{2}\cdot\partial_{\bm{\mathsf{X}}}J_{\Sigma}+\epsilon\,r^{-1}([u\,\bm{b}+r\,\bm{e}_{1}]\cdot\nabla\bm{b}\cdot\bm{e}_{1})\,\partial_{u}J_{\Sigma}+r^{-1}|\bm{B}|\partial_{r}J_{\Sigma}$
$+\epsilon\,r^{-1}\,(\bm{b}\cdot\nabla\times\bm{b})(u\partial_{r}J_{\Sigma}-r\partial_{u}J_{\Sigma})+\epsilon\,(\partial_{u}J_{\Sigma}\,\bm{b}+\partial_{r}J_{\Sigma}\,\bm{e}_{1})\cdot(\nabla\bm{e}_{1}\cdot\bm{e}_{2}).$ (13)
Proof.
We will perform the calculation for the system $\dot{\bm{v}}=\bm{v}\times\bm{B}$, $\dot{\bm{x}}=\bm{v}$. The result in the Theorem statement follows from this calculation after applying the substitutions $t\rightarrow t/\epsilon$ and $\bm{B}\rightarrow\bm{B}/\epsilon$.
[…]
$\{F,G\}=\partial_{\bm{x}}F\cdot\partial_{\bm{v}}G-\partial_{\bm{x}}G\cdot\partial_{\bm{v}}F+\bm{B}\cdot\partial_{\bm{v}}F\times\partial_{\bm{v}}G.$
In terms of the coordinates $(\bm{x},v_{\parallel},v_{\perp},\zeta)$ defined by $\bm{v}=v_{\parallel}\bm{b}+v_{\perp}(\cos\zeta\,\bm{e}_{1}+\sin\zeta\,\bm{e}_{2})$, the bracket is instead
$\{F,G\}$ $=\bm{b}\cdot(\partial_{\bm{x}}F\,\partial_{v_{\parallel}}G-\partial_{\bm{x}}G\,\partial_{v_{\parallel}}F)+\bm{a}\cdot(\partial_{\bm{x}}F\,\partial_{v_{\perp}}G-\partial_{\bm{x}}G\,\partial_{v_{\perp}}F)$
$+(v_{\perp}\,\bm{a}\cdot\nabla\bm{b}\cdot\bm{a}+v_{\parallel}\,\bm{b}\cdot\nabla\bm{b}\cdot\bm{a})(\partial_{v_{\parallel}}F\,\partial_{v_{\perp}}G-\partial_{v_{\parallel}}G\,\partial_{v_{\perp}}F)$
$+\bigg{(}N^{\bm{x}}\cdot\partial_{\bm{x}}F+N^{v_{\parallel}}\,\partial_{v_{\parallel}}F+N^{v_{\perp}}\,\partial_{v_{\perp}}F\bigg{)}\,\partial_{\zeta}G$
$-\bigg{(}N^{\bm{x}}\cdot\partial_{\bm{x}}G+N^{v_{\parallel}}\,\partial_{v_{\parallel}}G+N^{v_{\perp}}\,\partial_{v_{\perp}}G\bigg{)}\,\partial_{\zeta}F,$
where $\bm{a}=\cos\zeta\,\bm{e}_{1}+\sin\zeta\,\bm{e}_{2}$, $\bm{c}=\bm{a}\times\bm{b}=-\cos\zeta\,\bm{e}_{2}+\sin\zeta\,\bm{e}_{1}$, and the components of $N$ are given by
$N^{\bm{x}}$ $=-\frac{1}{v_{\perp}}\bm{c}$ (14)
$N^{v_{\parallel}}$ $=-\bm{c}\cdot\nabla\bm{b}\cdot\bm{a}-\frac{v_{\parallel}}{v_{\perp}}(\bm{b}\cdot\nabla\bm{b})\cdot\bm{c}+\bm{b}\cdot\nabla\bm{e}_{1}\cdot\bm{e}_{2}$ (15)
$N^{v_{\perp}}$ $=\frac{|\bm{B}|}{v_{\perp}}+\frac{v_{\parallel}}{v_{\perp}}\,\bm{b}\cdot\nabla\times\bm{b}+\bm{a}\cdot\nabla\bm{e}_{1}\cdot\bm{e}_{2}.$ (16)
Theorem 2 therefore implies that the Poisson bracket {,}$\{\cdot,\cdot\}_{\Sigma}$ on the guiding center Poincaré section is given by,
$\{f,g\}_{\Sigma}$ $=\bm{b}\cdot(\partial_{\bm{\mathsf{X}}}f\,\partial_{u}g-\partial_{\bm{\mathsf{X}}}g\,\partial_{u}f)+\bm{e}_{1}\cdot(\partial_{\bm{\mathsf{X}}}f\,\partial_{r}g-\partial_{\bm{\mathsf{X}}}g\,\partial_{r}f)$
$+(r\,\bm{e}_{1}\cdot\nabla\bm{b}\cdot\bm{e}_{1}+u\,\bm{b}\cdot\nabla\bm{b}\cdot\bm{e}_{1})(\partial_{u}f\,\partial_{r}g-\partial_{u}g\,\partial_{r}f)$
$+\bigg{(}\bm{b}\cdot(\partial_{\bm{\mathsf{X}}}J_{\Sigma}\,\partial_{u}f-\partial_{\bm{\mathsf{X}}}f\,\partial_{u}J_{\Sigma})+\bm{e}_{1}\cdot(\partial_{\bm{\mathsf{X}}}J_{\Sigma}\,\partial_{r}f-\partial_{\bm{\mathsf{X}}}f\,\partial_{r}J_{\Sigma})$
$+(r\,\bm{e}_{1}\cdot\nabla\bm{b}\cdot\bm{e}_{1}+u\,\bm{b}\cdot\nabla\bm{b}\cdot\bm{e}_{1})(\partial_{u}J_{\Sigma}\,\partial_{r}f-\partial_{u}f\,\partial_{r}J_{\Sigma})\bigg{)}\frac{N^{\bm{\mathsf{X}}}_{\Sigma}\cdot\partial_{\bm{\mathsf{X}}}g+N^{u}_{\Sigma}\,\partial_{u}g+N^{r}_{\Sigma}\,\partial_{r}g}{N^{\bm{\mathsf{X}}}_{\Sigma}\cdot\partial_{\bm{\mathsf{X}}}J_{\Sigma}+N^{u}_{\Sigma}\,\partial_{u}J_{\Sigma}+N^{r}_{\Sigma}\,\partial_{r}J_{\Sigma}}$
$-\bigg{(}\bm{b}\cdot(\partial_{\bm{\mathsf{X}}}J_{\Sigma}\,\partial_{u}g-\partial_{\bm{\mathsf{X}}}g\,\partial_{u}J_{\Sigma})+\bm{e}_{1}\cdot(\partial_{\bm{\mathsf{X}}}J_{\Sigma}\,\partial_{r}g-\partial_{\bm{\mathsf{X}}}g\,\partial_{r}J_{\Sigma})$
$+(r\,\bm{e}_{1}\cdot\nabla\bm{b}\cdot\bm{e}_{1}+u\,\bm{b}\cdot\nabla\bm{b}\cdot\bm{e}_{1})(\partial_{u}J_{\Sigma}\,\partial_{r}g-\partial_{u}g\,\partial_{r}J_{\Sigma})\bigg{)}\frac{N^{\bm{\mathsf{X}}}_{\Sigma}\cdot\partial_{\bm{\mathsf{X}}}f+N^{u}_{\Sigma}\,\partial_{u}f+N^{r}_{\Sigma}\,\partial_{r}f}{N^{\bm{\mathsf{X}}}_{\Sigma}\cdot\partial_{\bm{\mathsf{X}}}J_{\Sigma}+N^{u}_{\Sigma}\,\partial_{u}J_{\Sigma}+N^{r}_{\Sigma}\,\partial_{r}J_{\Sigma}},$
where
$N^{\bm{\mathsf{X}}}_{\Sigma}$ $=r^{-1}\bm{e}_{2}$
$N^{u}_{\Sigma}$ $=r^{-1}(u\bm{b}+r\bm{e}_{1})\cdot\nabla\bm{b}\cdot\bm{e}_{2}-\bm{b}\cdot\nabla\times\bm{b}+\bm{b}\cdot\nabla\bm{e}_{1}\cdot\bm{e}_{2}$
$N^{r}_{\Sigma}$ $=r^{-1}|\bm{B}|+r^{-1}u\,\bm{b}\cdot\nabla\times\bm{b}+\bm{e}_{1}\cdot\nabla\bm{e}_{1}\cdot\bm{e}_{2}.$
By Theorem 2, the footpoint flow field for the Lorentz force is given by Hamilton’s equations $\dot{\sigma}^{i}=\{\sigma^{i},H_{\Sigma}\}_{\Sigma}$. The above explicit expression for {,}$\{\cdot,\cdot\}_{\Sigma}$, together with $H_{\Sigma}=\frac{1}{2}u^{2}+\frac{1}{2}r^{2}$, therefore implies the following explicit expressions for the components of the footpoint flow field. The $\bm{X}$-component is
$\{\bm{\mathsf{X}},H_{\Sigma}\}_{\Sigma}$ $=u\,\bm{b}+r\,\bm{e}_{1}$
$+\bigg{(}-\partial_{u}J_{\Sigma}\,\bm{b}-\partial_{r}J_{\Sigma}\,\bm{e}_{1}\bigg{)}\frac{N^{u}_{\Sigma}\,u+N^{r}_{\Sigma}\,r}{N^{\bm{\mathsf{X}}}_{\Sigma}\cdot\partial_{\bm{\mathsf{X}}}J_{\Sigma}+N^{u}_{\Sigma}\,\partial_{u}J_{\Sigma}+N^{r}_{\Sigma}\,\partial_{r}J_{\Sigma}}$
$-\bigg{(}(u\,\bm{b}+r\,\bm{e}_{1})\cdot\partial_{\bm{\mathsf{X}}}J_{\Sigma}$
$+(r\,\bm{e}_{1}\cdot\nabla\bm{b}\cdot\bm{e}_{1}+u\,\bm{b}\cdot\nabla\bm{b}\cdot\bm{e}_{1})(\partial_{u}J_{\Sigma}\,r-u\,\partial_{r}J_{\Sigma})\bigg{)}\frac{N^{\bm{\mathsf{X}}}_{\Sigma}}{N^{\bm{\mathsf{X}}}_{\Sigma}\cdot\partial_{\bm{\mathsf{X}}}J_{\Sigma}+N^{u}_{\Sigma}\,\partial_{u}J_{\Sigma}+N^{r}_{\Sigma}\,\partial_{r}J_{\Sigma}},$
$=u\,\bm{b}+r\,\bm{e}_{1}$
$-\frac{(\partial_{u}J_{\Sigma}\,\bm{b}+\partial_{r}J_{\Sigma}\,\bm{e}_{1})(N^{u}_{\Sigma}\,u+N^{r}_{\Sigma}\,r)}{N^{\bm{\mathsf{X}}}_{\Sigma}\cdot\partial_{\bm{\mathsf{X}}}J_{\Sigma}+N^{u}_{\Sigma}\,\partial_{u}J_{\Sigma}+N^{r}_{\Sigma}\,\partial_{r}J_{\Sigma}}$
$-\frac{(u\,\bm{b}+r\,\bm{e}_{1})\cdot\partial_{\bm{\mathsf{X}}}J_{\Sigma}\,N^{\bm{\mathsf{X}}}_{\Sigma}}{N^{\bm{\mathsf{X}}}_{\Sigma}\cdot\partial_{\bm{\mathsf{X}}}J_{\Sigma}+N^{u}_{\Sigma}\,\partial_{u}J_{\Sigma}+N^{r}_{\Sigma}\,\partial_{r}J_{\Sigma}}$
$-\frac{(r\,\bm{e}_{1}\cdot\nabla\bm{b}\cdot\bm{e}_{1}+u\,\bm{b}\cdot\nabla\bm{b}\cdot\bm{e}_{1})(\partial_{u}J_{\Sigma}\,r-u\,\partial_{r}J_{\Sigma})N^{\bm{\mathsf{X}}}_{\Sigma}}{N^{\bm{\mathsf{X}}}_{\Sigma}\cdot\partial_{\bm{\mathsf{X}}}J_{\Sigma}+N^{u}_{\Sigma}\,\partial_{u}J_{\Sigma}+N^{r}_{\Sigma}\,\partial_{r}J_{\Sigma}}$
$=u\,\bm{b}+r\,\bm{e}_{1}$
$-\frac{\partial_{u}J_{\Sigma}(uN^{u}_{\Sigma}+r\,N^{r}_{\Sigma})}{N^{\bm{\mathsf{X}}}_{\Sigma}\cdot\partial_{\bm{\mathsf{X}}}J_{\Sigma}+N^{u}_{\Sigma}\,\partial_{u}J_{\Sigma}+N^{r}_{\Sigma}\,\partial_{r}J_{\Sigma}}\,\bm{b}$
$-\frac{\partial_{r}J_{\Sigma}(uN^{u}_{\Sigma}+r\,N^{r}_{\Sigma})}{N^{\bm{\mathsf{X}}}_{\Sigma}\cdot\partial_{\bm{\mathsf{X}}}J_{\Sigma}+N^{u}_{\Sigma}\,\partial_{u}J_{\Sigma}+N^{r}_{\Sigma}\,\partial_{r}J_{\Sigma}}\,\bm{e}_{1}$
$-\frac{(\frac{u}{r}\bm{b}+\bm{e}_{1})\cdot\partial_{\bm{\mathsf{X}}}J_{\Sigma}+(\bm{e}_{1}\cdot\nabla\bm{b}\cdot\bm{e}_{1}+\frac{u}{r}\,\bm{b}\cdot\nabla\bm{b}\cdot\bm{e}_{1})(r\,\partial_{u}J_{\Sigma}-u\,\partial_{r}J_{\Sigma})}{N^{\bm{\mathsf{X}}}_{\Sigma}\cdot\partial_{\bm{\mathsf{X}}}J_{\Sigma}+N^{u}_{\Sigma}\,\partial_{u}J_{\Sigma}+N^{r}_{\Sigma}\,\partial_{r}J_{\Sigma}}\,\bm{e}_{2}$
$=u\,\bm{b}+r\,\bm{e}_{1}$
$-\frac{\partial_{u}J_{\Sigma}(|\bm{B}|+u\,\bm{b}\cdot\nabla\times\bm{b}+u\,\bm{e}_{2}\cdot\nabla\bm{b}\cdot\bm{e}_{1}+\frac{u^{2}}{r}\bm{b}\cdot\nabla\bm{b}\cdot\bm{e}_{2}+(u\bm{b}+r\,\bm{e}_{1})\cdot\nabla\bm{e}_{1}\cdot\bm{e}_{2})}{N^{\bm{\mathsf{X}}}_{\Sigma}\cdot\partial_{\bm{\mathsf{X}}}J_{\Sigma}+N^{u}_{\Sigma}\,\partial_{u}J_{\Sigma}+N^{r}_{\Sigma}\,\partial_{r}J_{\Sigma}}\,\bm{b}$
$-\frac{\partial_{r}J_{\Sigma}(|\bm{B}|+u\,\bm{b}\cdot\nabla\times\bm{b}+u\,\bm{e}_{2}\cdot\nabla\bm{b}\cdot\bm{e}_{1}+\frac{u^{2}}{r}\bm{b}\cdot\nabla\bm{b}\cdot\bm{e}_{2}+(u\bm{b}+r\,\bm{e}_{1})\cdot\nabla\bm{e}_{1}\cdot\bm{e}_{2})}{N^{\bm{\mathsf{X}}}_{\Sigma}\cdot\partial_{\bm{\mathsf{X}}}J_{\Sigma}+N^{u}_{\Sigma}\,\partial_{u}J_{\Sigma}+N^{r}_{\Sigma}\,\partial_{r}J_{\Sigma}}\,\bm{e}_{1}$
$-\frac{(\frac{u}{r}\bm{b}+\bm{e}_{1})\cdot\partial_{\bm{\mathsf{X}}}J_{\Sigma}+(\bm{e}_{1}\cdot\nabla\bm{b}\cdot\bm{e}_{1}+\frac{u}{r}\,\bm{b}\cdot\nabla\bm{b}\cdot\bm{e}_{1})(r\,\partial_{u}J_{\Sigma}-u\,\partial_{r}J_{\Sigma})}{N^{\bm{\mathsf{X}}}_{\Sigma}\cdot\partial_{\bm{\mathsf{X}}}J_{\Sigma}+N^{u}_{\Sigma}\,\partial_{u}J_{\Sigma}+N^{r}_{\Sigma}\,\partial_{r}J_{\Sigma}}\,\bm{e}_{2},$
which reproduces Eqs. (8)-(10) in the Theorem statement. The $u$-component is
$\{u,H_{\Sigma}\}_{\Sigma}$ $=$
$+r\,(r\,\bm{e}_{1}\cdot\nabla\bm{b}\cdot\bm{e}_{1}+u\,\bm{b}\cdot\nabla\bm{b}\cdot\bm{e}_{1})$
$+\bigg{(}\bm{b}\cdot\partial_{\bm{\mathsf{X}}}J_{\Sigma}-(r\,\bm{e}_{1}\cdot\nabla\bm{b}\cdot\bm{e}_{1}+u\,\bm{b}\cdot\nabla\bm{b}\cdot\bm{e}_{1})\partial_{r}J_{\Sigma}\bigg{)}\frac{N^{u}_{\Sigma}\,u+N^{r}_{\Sigma}\,r}{N^{\bm{\mathsf{X}}}_{\Sigma}\cdot\partial_{\bm{\mathsf{X}}}J_{\Sigma}+N^{u}_{\Sigma}\,\partial_{u}J_{\Sigma}+N^{r}_{\Sigma}\,\partial_{r}J_{\Sigma}}$
$-\bigg{(}(u\,\bm{b}+r\,\bm{e}_{1})\cdot\partial_{\bm{\mathsf{X}}}J_{\Sigma}$
$+(r\,\bm{e}_{1}\cdot\nabla\bm{b}\cdot\bm{e}_{1}+u\,\bm{b}\cdot\nabla\bm{b}\cdot\bm{e}_{1})(\partial_{u}J_{\Sigma}\,r-u\,\partial_{r}J_{\Sigma})\bigg{)}\frac{N^{u}_{\Sigma}}{N^{\bm{\mathsf{X}}}_{\Sigma}\cdot\partial_{\bm{\mathsf{X}}}J_{\Sigma}+N^{u}_{\Sigma}\,\partial_{u}J_{\Sigma}+N^{r}_{\Sigma}\,\partial_{r}J_{\Sigma}}$
$=\frac{r\,(r\bm{e}_{1}\cdot\nabla\bm{b}\cdot\bm{e}_{1}+u\bm{b}\cdot\nabla\bm{b}\cdot\bm{e}_{1})\,N^{\bm{\mathsf{X}}}_{\Sigma}\cdot\partial_{\bm{\mathsf{X}}}J_{\Sigma}}{D}$
$+\frac{r(N^{r}_{\Sigma}\,\bm{b}-N^{u}_{\Sigma}\,\bm{e}_{1})\cdot\partial_{\bm{\mathsf{X}}}J_{\Sigma}}{D}$
$=\frac{(|\bm{B}|+u\bm{b}\cdot\nabla\times\bm{b}+r\,\bm{e}_{1}\cdot\nabla\bm{e}_{1}\cdot\bm{e}_{2})}{D}\bm{b}\cdot\partial_{\bm{\mathsf{X}}}J_{\Sigma}$
$-\frac{(r\bm{e}_{1}\cdot\nabla\bm{b}\cdot\bm{e}_{2}-r\bm{b}\cdot\nabla\times\bm{b}+u\bm{b}\cdot\nabla\bm{b}\cdot\bm{e}_{2}+r\bm{b}\cdot\nabla\bm{e}_{1}\cdot\bm{e}_{2})}{D}\bm{e}_{1}\cdot\partial_{\bm{\mathsf{X}}}J_{\Sigma}$
$+\frac{(r\,\bm{e}_{1}\cdot\nabla\bm{b}\cdot\bm{e}_{1}+u\bm{b}\cdot\nabla\bm{b}\cdot\bm{e}_{1})}{D}\bm{e}_{2}\cdot\partial_{\bm{\mathsf{X}}}J_{\Sigma}$
$=\frac{\bm{B}+(\bm{b}\cdot\nabla\times\bm{b})[u\,\bm{b}+r\,\bm{e}_{1}]+\bm{b}\times([u\bm{b}+r\bm{e}_{1}]\cdot\nabla\bm{b})-r\,\bm{e}_{2}\times(\nabla\bm{e}_{1}\cdot\bm{e}_{2})}{D}\cdot\partial_{\bm{\mathsf{X}}}J_{\Sigma},$
which reproduces Eq. (11). Finally, the $r$-component is
$\{r,H_{\Sigma}\}_{\Sigma}$ $=-u(r\,\bm{e}_{1}\cdot\nabla\bm{b}\cdot\bm{e}_{1}+u\,\bm{b}\cdot\nabla\bm{b}\cdot\bm{e}_{1})$
$+\bigg{(}\bm{e}_{1}\cdot(\partial_{\bm{\mathsf{X}}}J_{\Sigma})$
$+(r\,\bm{e}_{1}\cdot\nabla\bm{b}\cdot\bm{e}_{1}+u\,\bm{b}\cdot\nabla\bm{b}\cdot\bm{e}_{1})(\partial_{u}J_{\Sigma})\bigg{)}\frac{N^{u}_{\Sigma}\,u+N^{r}_{\Sigma}\,r}{N^{\bm{\mathsf{X}}}_{\Sigma}\cdot\partial_{\bm{\mathsf{X}}}J_{\Sigma}+N^{u}_{\Sigma}\,\partial_{u}J_{\Sigma}+N^{r}_{\Sigma}\,\partial_{r}J_{\Sigma}}$
$-\bigg{(}(u\,\bm{b}+r\,\bm{e}_{1})\cdot\partial_{\bm{\mathsf{X}}}J_{\Sigma}$
$+(r\,\bm{e}_{1}\cdot\nabla\bm{b}\cdot\bm{e}_{1}+u\,\bm{b}\cdot\nabla\bm{b}\cdot\bm{e}_{1})(\partial_{u}J_{\Sigma}\,r-u\,\partial_{r}J_{\Sigma})\bigg{)}\frac{N^{r}_{\Sigma}}{N^{\bm{\mathsf{X}}}_{\Sigma}\cdot\partial_{\bm{\mathsf{X}}}J_{\Sigma}+N^{u}_{\Sigma}\,\partial_{u}J_{\Sigma}+N^{r}_{\Sigma}\,\partial_{r}J_{\Sigma}}$
$=-\frac{u\,(r\,\bm{e}_{1}\cdot\nabla\bm{b}\cdot\bm{e}_{1}+u\,\bm{b}\cdot\nabla\bm{b}\cdot\bm{e}_{1})\,N^{\bm{\mathsf{X}}}_{\Sigma}}{D}\cdot\partial_{\bm{\mathsf{X}}}J_{\Sigma}$
$+\frac{u\,N^{u}_{\Sigma}\,\bm{e}_{1}-u\,N^{r}_{\Sigma}\,\bm{b}}{D}\cdot\partial_{\bm{\mathsf{X}}}J_{\Sigma}$
$=\frac{-\frac{u}{r}\bm{b}\times([u\bm{b}+r\bm{e}_{1}]\cdot\nabla\bm{b})-\frac{u}{r}[u\bm{b}+r\bm{e}_{1}](\bm{b}\cdot\nabla\times\bm{b})+u\bm{e}_{2}\times(\nabla\bm{e}_{1}\cdot\bm{e}_{2})}{D}\cdot\partial_{\bm{\mathsf{X}}}J_{\Sigma},$
which reproduced Eq. (12) from the Theorem statement.
I.2 Derivative estimation
This Section describes the method we used to estimate “ground truth” time derivatives of (presumptive) continuous-time trajectories on $\Sigma$ that interpolate iterates of the Poincaré map. To provide good input data to estimate $\mathcal{J}$, we need to be confident that we can both trust that a point $(x,y,r)$ is integrable and that we have a good estimate of the derivative $(\dot{x},\dot{y})$. For both of these tasks, we leverage the Birkhoff Reduced Rank Extrapolation (Birkhoff RRE) algorithm Ruth:2024 implemented in the SymplecticMapTools.jl Julia package. Using a single trajectory from the Poincaré map, here denoted $F$, Birkhoff RRE first classifies that trajectory as an invariant circle, an island, or chaos. Then, assuming the trajectory is an invariant circle with Diophantine rotation number $\omega$, it returns an approximate parameterization of an invariant circle $z:\mathbb{T}\to\Sigma$ and $\omega$ so that $F(z(\theta))=z(\theta+\omega)$. Clearly, the adiabatic invariant $\mathcal{J}$ should be constant on $z$, meaning that $\dot{z}\cdot\nabla\mathcal{J}=0$, giving the correct direction of the derivatives of $\dot{z}$ in 2D.
To obtain the magnitude of of the derivatives, we first observe that the dynamics predicted by the non-perturbative guiding center equations of motion on the invariant circle will have the form $(\dot{x},\dot{y})=({dz}/{d\theta})\dot{\theta}$. To estimate $\dot{\theta}$, we note it must satisfy a differential equation of the form
$\dot{\theta}=f(\theta),\qquad\theta(t+T(\theta(t)))=\theta(t)+\omega,$
where $T(\theta(t))$ is the gyroperiod at the point $z(\theta(t))$. We can lift and invert the dynamics to find
$\frac{dt}{d\theta}=g(t),\qquad t(\theta+\omega)=t(\theta)+T(\theta).$
For Diophantine $\omega$ and smooth enough $T$, we can solve the right expression for $t$ to find
$t(\theta)=c+t_{0}\theta+\sideset{}{{}^{\prime}}{\sum}_{n=-N}^{N}t_{n}e^{in\theta},$
where $c$ is a constant determined by initial conditions and
[…]
Using a discrete Fourier transform to compute the above quantities and $\dot{\theta}=({dt}/{d\theta})^{-1}$, we have an approximation of $(\dot{x},\dot{y})$.